# CMB parity asymmetry from unitary quantum gravitational physics

### v2

@@ Abstract
Longstanding anomalies in the Cosmic Microwave Background (CMB), including the low quadrupole moment and hemispherical power asymmetry, have recently been linked to an underlying parity asymmetry. We show here how this parity asymmetry naturally arises within a quantum framework that explicitly incorporates PT symmetry—parity (P) and time-reversal (T). This additional symmetry restores unitarity in quantum field theory (QFT) in curved spacetime. When applied to inflationary quantum fluctuations, this extended symmetry predicts parity asymmetry as a natural consequence of cosmic expansion, which inherently breaks time-reversal symmetry. Observational data strongly favor this unitary approach, with a likelihood ratio exceeding 650 times that of predictions from the standard inflationary framework. These results provide compelling evidence for the quantum gravitational origins of CMB parity asymmetry.
@@ I Introduction
Understanding and testing the interplay between gravity and quantum mechanics (QM) is the most important key to unraveling the new physics beyond the standard model of particle physics and Einstein’s general relativity (GR). The merge of GR and QM is an essential physics that can be unlocked at the length scales of gravitational horizons. For this, we perhaps need an in-depth understanding of how one can resolve an age-old conundrum in the concept of time between the GR and QM. The success of Quantum field theory (QFT) in Minkowski spacetime (Coleman (2018)) already teaches us several lessons in this regard. Indeed, QFT is the culmination of special relativity (SR) and QM, where the concept of time already enters into the conflict. In SR, the space and time are on equal footing, but in QM, spatial coordinates are elevated to operators while time is just a parameter due to its anti-unitary nature under the discrete transformation ($\mathcal{T}:t\to-t$). The SR and QM are merged by imposing the commutativity of operators corresponding to space-like distances. The underlying reason is that SR forbids any communication happening in space-like distances, and we respect it and achieve it by formulating new mathematical rules via second quantization. This is exactly the lesson one must keep in mind to understand quantum fields in curved spacetime. That is, we take lessons from classical physics and comply with them according to the rules of quantum physics. The quantum field theory in curved spacetime (QFTCS) is the first step in seeing GR and QM acting together besides the renormalizable ultra-violet complete quantum gravity that may be required at the Planck scales. Furthermore, QFTCS is a necessity in the context of understanding inflationary quantum fluctuations (Mukhanov and Chibisov (1981); Sasaki (1986); Albrecht et al. (1994)).
@@
In this paper, we address crucial questions related to the meaning of time reversal and parity inversions in the context of inflationary background spacetime to place quantum fields on it and uncover their true nature in the cosmic microwave background (CMB). We explain how direct-sum quantum field theory (DQFT), Kumar and Marto (2025, 2024); Gaztañaga et al. (2024)) is a necessity to solve the unitarity problem of QFTCS. The question of unitarity in QFTCS precedes the question of a completely renormalizable quantum gravity theory of Planck scales. In contrast to the widespread schools of thought that argue unitarity in QFTCS needs to be given up within GR and QM, we propose a new fundamental understanding of this issue and a new observational test using the parity asymmetry in the CMB. We find that the theory of inflationary quantum fluctuations with DQFT (which we call shortly DSI, which stands for ”Direct-Sum inflation”) fits up to 650 times better than the Standard Inflation (SI). This paper ultimately presents the picture of how a new theory of QFTCS emerged from fundamental questions of quantum gravitational physics that can potentially explain the observed CMB large-scale anomalies. The new results in this paper are further complemented by our more extended companion paper (Gaztañaga and Kumar (2024)).
@@ II DQFT, in a nutshell ($\hbar=c=1$)
DQFT formulation emerges from a simple reconstruction of a quantum state or a quantum field operator through the discrete (a)symmetries (parity $\mathcal{P}$ and time reversal $\mathcal{T}$) of the background spacetime. Here, we summarize the conceptual understanding of DQFT and direct the reader to Gaztañaga and Kumar (2024); Kumar and Marto (2025, 2024); Gaztañaga et al. (2024) for more details. Construction of QFT in Minkowski spacetime requires the definition of positive energy state and the arrow of time Donoghue and Menezes (2019). Here, we build quantum theory with two arrows of time using the direct-sum Schrödinger equation
@@
where $\mathbb{H}=\mathbb{H}_{+}\oplus\mathbb{H}_{-}$ is the time-independent Hamiltonian operator (which is Hermitian) that is $\mathcal{P}\mathcal{T}$ symmetric. The quantum state $|\Psi\rangle$ is the direct-sum of two components
@@
where $|\Psi_{+}\rangle,\,|\Psi_{-}\rangle$ are the two components of the same state at parity conjugate points. The component $|\Psi_{+}\rangle$ is a positive energy state that evolves forward in time as $|\Psi_{+}\rangle_{t_{p}}=e^{-i\mathbb{E}t_{p}}|\Psi_{+}\rangle_{0}$ where the parametric time arrow $t_{p}:-\infty\to\infty$ while $|\Psi_{-}\rangle$ can also be seen as a positive energy state that evolves back in time as $|\Psi_{-}\rangle_{t_{p}}=e^{i\mathbb{E}t_{p}}|\Psi_{-}\rangle_{0}$ with the parametric time convention $t_{p}:+\infty\to-\infty$. One can notice here the purely antiunitary character of time associated with the replacement $i\to-i$ (Donoghue and Menezes (2019)). The direct-sum Schrödinger equation does not have any arrow of time associated with it compared to the conventional Schrödinger equation. The direct-sum splitting of the quantum state as in Eq.2 implies the Hilbert space ($\mathcal{H}$) splitting into a direct-sum of the two ($\mathcal{H}_{\pm}$) corresponding to parity conjugate points in position space: $\mathcal{H}=\mathcal{H}_{+}\oplus\mathcal{H}_{-}$. Here, the Hilbert spaces $\mathcal{H}_{\pm}$ are called the geometric superselection sectors (Gaztañaga et al. (2024); Kumar and Marto (2025)). The position and momentum commutation relations now become double associated with parity conjugate position and momentum space operators denoted by subscripts +,+,\,-+ , -, and their relations are
$\hat{p}_{+}=-i\frac{\partial}{\partial x_{+}}$ ; $\Big{[}\hat{x}_{+},\,\hat{p}_{+}\Big{]}=-\Big{[}\hat{x}_{-},\,\hat{p}_{-}\Big{]}=i\,$ (3)
$\hat{p}_{-}=+i\frac{\partial}{\partial x_{-}}$ ; $[\hat{x}_{+},\,\hat{x}_{-}]=[\hat{p}_{+},\,\hat{p}_{-}]=[\hat{x}_{\pm},\,\hat{p}_{\mp}]=0\,.$
@@
This framework only makes the wavefunction of the quantum system explicitly $\mathcal{P}\mathcal{T}$ symmetric that resonates with the assumed $\mathcal{P}\mathcal{T}$ symmetry of the physical system, i.e. $[\mathbb{H},\,\mathcal{P}\mathcal{T}]=0$ and the results of standard quantum mechanics remain the same as shown with the example of harmonic oscillator worked out in Gaztañaga and Kumar (2024) (see also further discussions and Fig.1 in Kumar and Marto (2025)). The DQFT in Minkowski spacetime ($ds^{2}=-dt_{m}^{2}+d\textbf{x}^{2}$ which is $\mathcal{P}\mathcal{T}$ symmetric) is just built on the direct-sum Schrödinger equation, following how the standard QFT is built on the definition of positive energy state and the causality condition (i.e., commutativity of operators for spacelike distances). See Appendix A for more details.
@@ III DQFT in de Sitter & Unitarity ($G=\hbar=c=1$)
The de Sitter (dS) spacetime is a perfect testing ground for exploring the theory of QFTCS because of its maximally symmetric aspect, and it is also relevant for unlocking the nature of inflationary quantum fluctuations. Furthermore, a dS spacetime has the closest resemblance with black holes, where the most important conundrums in quantum physics emerge such as the loss of unitarity and the information-loss paradox. We start with the dS metric in the flat Friedman-Lemaître-Robertson-Walker (FLRW) coordinates
@@
where the scale factor $a=e^{Ht}$ is the clock that determines the expansion of the Universe, $H=\frac{\dot{a}}{a}$ is the Hubble parameter and $\tau=\int\frac{dt}{a}=\frac{1}{aH}$ is the conformal time. If one has to place quantum theory in dS spacetime Eq.4 the first and foremost things to be understood are the parity and time reversal operations. The usual conception of understanding the expansion of the Universe is to restrict ourselves to $\tau<0$. But the metric carries a symmetry
@@
Restricting to $\tau<0$ before quantization of a classical field (as is usually followed in standard QFTCS) means throwing away the symmetry in Eq.5 by hand. The expanding Universe can be realized in two ways:
@@
This would explain why in the Wheeler-de Witt equation the cosmic time does not explicitly appear, rather the scale factor is the clock that runs with expanding Universe (Kiefer (2007)). Note that $H\to-H$ does not change any properties of the dS space at all. For example, curvature invariants such as the Ricci scalar $R=12H^{2}$ remain the same under a discrete sign flip of H. Eq.6 illustrates two possible time realizations for an expanding universe. By convention, choosing $\tau>0$ or $\tau<0$ results in loss of information beyond the horizon. In his famous monograph of 1956, Schrodinger Erwin Schrodinger (1956) rejected the idea of two universes because it would lead to a conundrum of ignorance of information beyond the horizon, which in modern language means pure states evolving into mixed states which leads the violation of unitarity and information-loss (Gibbons and Hawking (1977); Parikh et al. (2003)).
Similar to the Minkowski space (see Appendix A), the quantum fields in the dS space are described through a direct-sum vacuum $|0_{T}\rangle_{\rm dS}=|0_{+}\rangle_{\rm dS}\oplus|0_{-}\rangle_{\rm dS}$ corresponding to the direct-sum Fock space $\mathcal{F}_{\rm dS}=\mathcal{F}_{\rm dS+}\oplus\mathcal{F}_{\rm dS-}$. This implies that everywhere in dS space, a quantum field is tied by time forward and backward evolutions at the parity conjugate points, see Appendix B. At any moment of dS expansion, the parity-conjugate points on the comoving horizon $r_{H}=|\frac{1}{aH}|$ are space-like separated for an imaginary classical observer as shown in Fig. 1.
@@
Note that by DQFT construction, the direct-sum quantum states $|\phi_{\pm}\rangle=\hat{d}_{(\pm)\,\textbf{k}}^{\dagger}|0_{\pm}\rangle_{\rm dS}$ that follow from Eq.26 are the opposite time evolving positive energy states at the parity conjugate points and this holds for any imaginary classical observer. Furthermore, in DQFT the causality and locality are preserved too. A consequence of this is that any state that disappears beyond the horizon happens to be the complimentary direct-sum (component) state that reappears at the antipodal point. This is because all the imaginary classical observers, including those on the horizons of each other, should observe the same direct-sum quantum state $|\phi_{+}\rangle\oplus|\phi_{-}\rangle$.
@@
where $|\phi_{m}\rangle$ are the eigenvectors, i.e. for any observable (an operator) $\hat{O}$ we have: $\hat{O}|\phi_{m}\rangle=\lambda_{m}|\phi_{m}\rangle$ where $\lambda_{m}$ are the corresponding eigenvalues. Each $\lambda_{m}$ will be observed in $|\psi\rangle$ with a probability: $P(\lambda_{m})=|d_{m}|^{2}$. Consider now the (deSitter) Hubble horizon $r_{H}=|\frac{1}{aH}|$ around an observer (as illustrated in Fig.1). In the standard approach, information is lost (a pure state becomes a mix state) because communication can not happen for space-like distances (i.e. outside the horizon) as was shown by Gibbons and Hawking in 1977 Gibbons and Hawking (1977), in analogy to what happens in Hawking radiation and the black hole information loss paradox 1974Natur.248...30H; Kumar and Marto (2024). This also implies the lost of unitarity in QFT on spacetimes with horizons. In the DQFT approach we have:
@@
where $|\phi_{Am}^{\pm}\rangle$ and $|\phi_{Bn}^{\pm}\rangle$ are the eigenvectors corresponding to each $\mathcal{P}\mathcal{T}$ separated Hilbert spaces, which remain within the horizon because no information is exchanged across $\mathcal{P}\mathcal{T}$ separated spaces. All information about the entanglement pairs is mapped within the horizon; thus, the pure state will remain pure. In other words, DQFT forbids any entanglement beyond the gravitational horizons. DQFT constructs a direct-sum Hilbert space with the interior and exterior separated by a geometric superselection rule (which is similar to the concept of superselection sectors in QFT Wick et al. (1952); nLab authors (2023)). Thus, there would be an observer complementarity (see Fig.3 in Kumar and Marto (2025)). We can also understand the same from the point of view of dS space in static coordinates, which reads as

### v4

@@ Abstract
Longstanding anomalies in the Cosmic Microwave Background (CMB), including the low quadrupole moment and hemispherical power asymmetry, have recently been linked to an underlying parity asymmetry. We show here how this parity asymmetry naturally arises within a quantum framework that explicitly incorporates the construction of a geometric quantum vacuum based on parity ($\mathcal{P}$) and time-reversal ($\mathcal{T}$) transformations. This elegant framework restores unitarity in quantum field theory in curved spacetime (QFTCS). When applied to inflationary quantum fluctuations, this unitary QFTCS formalism predicts parity asymmetry as a natural consequence of cosmic expansion, which inherently breaks time-reversal symmetry. Observational data strongly favor this unitary QFTCS approach, with a Bayes factor, the ratio of marginal likelihoods associated with the model given the data $p\left(M|D\right)$, exceeding 650 times that of predictions from the standard inflationary framework. This Bayesian approach contrasts with the standard practice in the CMB community, which evaluates $p\left(D|M\right)$, the likelihood of the data under the model, which undermines the importance of $-\ell$ physics. Our results, for the first time, provide compelling evidence for the quantum gravitational origins of CMB parity asymmetry on large scales.
@@ I Introduction
Understanding and testing the interplay between gravity and quantum mechanics (QM) is the most important key to unraveling the new physics beyond the standard model of particle physics and Einstein’s general relativity (GR). The merger of GR and QM is an essential physics that can be unlocked at the length scales of gravitational horizons. For this, we perhaps need an in-depth understanding of how one can resolve an age-old conundrum in the concept of time between the GR and QM. The success of Quantum field theory (QFT) in Minkowski spacetime (Coleman (2018)) already teaches us several lessons in this regard. Indeed, QFT is the culmination of special relativity (SR) and QM, where the concept of time already enters into the conflict. In SR, the space and time are on equal footing, but in QM, spatial coordinates are elevated to operators while time is just a parameter due to its anti-unitary nature under the discrete transformation ($\mathcal{T}:t\to-t$). The SR and QM are merged by imposing the commutativity of operators corresponding to space-like distances. The underlying reason is that SR forbids any communication happening in space-like distances, and we respect it and achieve it by formulating new mathematical rules via second quantization. This is exactly the lesson one must keep in mind to understand quantum fields in curved spacetime. That is, we take lessons from classical physics and comply with them according to the rules of quantum physics. The quantum field theory in curved spacetime (QFTCS) is the first step in seeing GR and QM acting together, besides the renormalizable ultra-violet complete quantum gravity that may be required at the Planck scales. Furthermore, QFTCS is a necessity in the context of understanding inflationary quantum fluctuations (Mukhanov and Chibisov (1981); Sasaki (1986); Albrecht et al. (1994)). The longstanding problem of QFTCS has been the loss of unitarity and the information paradox, which was first formulated in the context of black holes by Hawking’s seminal works Hawking (1974, 1975) and later extended to de Sitter space in the paper of Gibbons and Hawking Gibbons and Hawking (1977). The problem of unitarity emerges with the way quantization is usually done in curved spacetime, leading to the phenomenon of pure states evolving into mixed states. It means, an observer bounded by a spacetime horizon loses access to the states beyond the horizon. This problem first identified by Schrödinger in 1956 in the context of de Sitter space Erwin Schrodinger (1956). Despite some attempts in the setups of de Sitter holography Parikh et al. (2003); Parikh and Verlinde (2005) which attempted to save unitarity, multiple investigations in recent decades have largely focused on accepting unitarity loss and proposing entangled pairs of universes and parallel spacetimes or restoring unitarity from ad-hoc yet unknown frameworks of quantum gravity, See Almheiri et al. (2021); Hartman et al. (2020); Shaghoulian and Susskind (2022); Balasubramanian et al. (2001, 2022); Shaghoulian (2022); Giddings (2022) and references therein. Motivated from ’t Hooft, Schrödinger and Einstein-Rosen’s seminal works ’t Hooft (2016); Erwin Schrodinger (1956); Einstein and Rosen (1935) which insisted on the role of parity and time reversal operations, the recent proposal of direct-sum quantum field theory (DQFT) restores the unitarity in curved spacetime Kumar and Marto (2025, 2024a); Gaztañaga et al. (2024); Kumar and Marto (2024b, c) along with the potential resolution to the information-loss paradox.
@@
In this paper, we address crucial questions related to the meaning of time reversal and parity inversions in the context of inflationary quantum fluctuations and uncover their true nature in the cosmic microwave background (CMB). We explain how DQFT Kumar and Marto (2025, 2024a); Gaztañaga et al. (2024)) is a necessity to solve the unitarity problem of QFTCS. The question of unitarity in QFTCS precedes the question of a completely renormalizable quantum gravity theory of Planck scales. In contrast to the widespread schools of thought that argue unitarity in QFTCS needs to be given up within GR and QM, we propose a new fundamental understanding of this issue and a new observational test using the parity asymmetry in the CMB. We find that the theory of inflationary quantum fluctuations with DQFT (which we call shortly DSI, which stands for ”Direct-Sum inflation”) fits up to 650 times better than the Standard Inflation (SI). This paper ultimately presents the picture of how a new theory of QFTCS emerged from fundamental questions of quantum gravitational physics that can potentially explain the observed CMB (large-scale) anomalies. The new results in this paper are further complemented by our more extended companion paper (Gaztañaga and Kumar (2024)).
The paper is organized as follows: In Sec . II we summarize the formulation of DQFT, which is built on a new understanding of quantum theory with two arrows of time along with parity-based geometric superselection rules. In Sec. III, we discuss the construction of DQFT in de Sitter spacetime and demonstrate how we achieve unitarity in this setup. In Sec . IV, we present the application of DQFT to inflationary quantum fluctuations (i.e., DSI) and provide the strong statistical evidence for it over the standard framework of inflation (SI) using the latest CMB data from Planck. In Sec . V we conclude the important elements of our investigation, which expose the role of gravity and quantum mechanics in the early universe and the fresh perspective we bring to the subject with compelling observational evidence. In Appendices, A, B and C, we provide further essential details of direct-sum QM, DQFT in Miknowski and de Sitter spacetimes.
@@ II DQFT, in a nutshell ($\hbar=c=1$)
DQFT formulation emerges from a simple reconstruction of a quantum state or a quantum field operator through the discrete (a)symmetries (parity $\mathcal{P}$ and time reversal $\mathcal{T}$) of the background spacetime. Here, we summarize the conceptual understanding of DQFT and direct the reader to Gaztañaga and Kumar (2024); Kumar and Marto (2025, 2024a); Gaztañaga et al. (2024); Kumar and Marto (2024d, b, c) for more details. Construction of QFT in Minkowski spacetime requires the definition of a positive energy state and the arrow of time Donoghue and Menezes (2019). Here, we build quantum theory with two arrows of time using the direct-sum Schrödinger equation
@@
where
^$\hat{\mathbb{H}}\left(\hat{x},\,\hat{p}\right)=\hat{\mathbb{H}}_{+}\left(\hat{x}_{+},\,\hat{p}_{+}\right)\oplus\hat{\mathbb{H}}_{-}\left(\hat{x}_{-},\,\hat{p}_{-}\right)=\begin{pmatrix}\mathbb{H}_{+}&&0\\ 0&&\mathbb{H}_{-}\end{pmatrix}$ (2)
it is the time-independent Hamiltonian operator (which is Hermitian) which we assume here to be $\mathcal{P}\mathcal{T}$ symmetric for simplicity.  [^1] It is important to note that our construction is not related to $\mathcal{P}\mathcal{T}$ symmetric QM where systems with non-Hermitian Hamiltonians are studied Bender et al. (1999). In contrast, we work here with Hermitian Hamiltonian operators. Here, the position operator is split into direct-sum of two components ($\hat{x}_{\pm}$) as
$\hat{x}=\hat{x}_{+}\oplus\hat{x}_{-}=\begin{pmatrix}\hat{x}_{+}&&0\\ 0&&\hat{x}_{-}\end{pmatrix}$ (3)
which have parity-conjugate eigenvalues when they act on the states $|\Psi_{\pm}\rangle$. Similarly, the momentum operator is split as
$\hat{p}=\hat{p}_{+}\oplus\hat{p}_{-}=\begin{pmatrix}\hat{p}_{+}&&0\\ 0&&\hat{p}_{-}\end{pmatrix}$ (4)
with $\hat{p}_{\pm}$ as the momentum operators associated with the states $|\Psi_{\pm}\rangle$ that correspond to Hilbert spaces representing parity conjugate regions of physical space as explained further below. The quantum state $|\Psi\rangle$ in Eq. 1 is expressed as the direct-sum of two components
@@
where $|\Psi_{+}\rangle,\,|\Psi_{-}\rangle$ are the two components of the same state at parity conjugate points. The component $|\Psi_{+}\rangle$ is a positive energy state ($\mathcal{E}>0$) that evolves forward in time as $|\Psi_{+}\rangle_{t_{p}}=e^{-i\mathcal{E}t_{p}}|\Psi_{+}\rangle_{0}$ where the parametric time ($t_{p}$) arrow $t_{p}:-\infty\to\infty$ while $|\Psi_{-}\rangle$ can also be seen as a positive energy state that evolves back in time as $|\Psi_{-}\rangle_{t_{p}}=e^{i\mathcal{E}t_{p}}|\Psi_{-}\rangle_{0}$ with the parametric time convention $t_{p}:+\infty\to-\infty$. One can notice here the purely antiunitary character of time associated with the replacement $i\to-i$ (Donoghue and Menezes (2019)). Note that the minus sign in front of $\mathbb{H}_{-}$ in Eq. 1 represents the opposite arrow of time. Thus, the direct-sum Schrödinger equation does not have any arrow of time associated with it compared to the conventional Schrödinger equation. To be explicit, we can rewrite Eq. 1 as
$\begin{pmatrix}i\frac{\partial}{\partial_{t_{p}}}|\Psi_{+}\rangle\\ -i\frac{\partial}{\partial_{t_{p}}}|\Psi_{-}\rangle\end{pmatrix}=\begin{pmatrix}\hat{\mathbb{H}}_{+}&&0\\ 0&&\hat{\mathbb{H}}_{-}\end{pmatrix}\begin{pmatrix}|\Psi_{+}\rangle\\ |\Psi_{-}\rangle\end{pmatrix}$ (6)
The direct-sum splitting of the quantum state as in Eq. 5 implies the Hilbert space ($\mathcal{H}$) splitting into a direct-sum of the two ($\mathcal{H}_{\pm}$) corresponding to parity conjugate points in position space: $\mathcal{H}=\mathcal{H}_{+}\oplus\mathcal{H}_{-}$. Here, the Hilbert spaces $\mathcal{H}_{\pm}$ are called the geometric superselection sectors (Gaztañaga et al. (2024); Kumar and Marto (2025)). The position and momentum commutation relations now become double associated with parity conjugate position and momentum space operators denoted by subscripts +,+,\,-+ , -, and their relations are
$\Big{[}\hat{x},\,\hat{p}\Big{]}=i\implies\begin{cases}\Big{[}\hat{x}_{+},\,\hat{p}_{+}\Big{]}=-\Big{[}\hat{x}_{-},\,\hat{p}_{-}\Big{]}=i\,\\ \par[\hat{x}_{+},\,\hat{x}_{-}]=[\hat{p}_{+},\,\hat{p}_{-}]=[\hat{x}_{\pm},\,\hat{p}_{\mp}]=0\end{cases},\quad\hat{p}_{\pm}=\mp i\frac{\partial}{\partial x_{\pm}},\quad{\begin{aligned} x_{+}&=x\in(0,\infty]\\ x_{-}&=x\in[-\infty,0)\\ \end{aligned}}$
The concept of wavefunction in the direct-sum QM becomes
$\Psi(x)=\frac{1}{\sqrt{2}}\begin{pmatrix}\langle x_{+}|\quad\langle x_{-}|\end{pmatrix}\begin{pmatrix}|\Psi_{+}\rangle_{0}e^{-i\mathcal{E}t_{p}}\\ |\Psi_{-}\rangle_{0}e^{i\mathcal{E}t_{p}}\end{pmatrix}\implies\begin{cases}\frac{1}{\sqrt{2}}\Psi_{+}\left(x_{+}\right)e^{-i\mathcal{E}t_{p}},\quad x_{+}=x\gtrsim 0\\ \frac{1}{\sqrt{2}}\Psi_{-}\left(x_{-}\right)e^{i\mathcal{E}t_{p}},\quad x_{-}=x\lesssim 0\,.\end{cases}$ (7)
@@
This framework makes the wavefunction of the quantum system explicitly $\mathcal{P}\mathcal{T}$ symmetric that resonates with the assumed $\mathcal{P}\mathcal{T}$ symmetry of the physical system, i.e. $[\mathbb{H},\,\mathcal{P}\mathcal{T}]=0$ and the results of standard quantum mechanics remain the same as shown with the example of harmonic oscillator worked out in Gaztañaga and Kumar (2024) (see also further discussions and Fig. 1 in Kumar and Marto (2025, 2024b)). The DQFT in Minkowski spacetime ($ds^{2}=-dt_{m}^{2}+d\textbf{x}^{2}$ ($t_{m}$ being the time coordinate) which is $\mathcal{P}\mathcal{T}$ symmetric i.e., $t_{m}\to-t_{m}$, $\to-\textbf{x}$ x) is just built on the direct-sum Schrödinger equation Eq. 1, following how the standard QFT is built on the definition of positive energy state and the causality condition (i.e., commutativity of operators for spacelike distances). See Appendix. A and B for more details.
@@ III DQFT in de Sitter & Unitarity ($G=\hbar=c=1$)
The de Sitter (dS) spacetime is a perfect testing ground for exploring the theory of QFTCS because of its maximally symmetric aspect, and it is also relevant for unlocking the nature of inflationary quantum fluctuations. Furthermore, dS spacetime has the closest resemblance to black holes, where the most important conundrums in quantum physics emerge, such as the loss of unitarity and the information-loss paradox Gibbons and Hawking (1977). We start with the dS metric in the flat Friedman-Lemaître-Robertson-Walker (FLRW) coordinates
@@
where the scale factor $a=e^{Ht}$ is the clock that determines the expansion of the Universe, $H=\frac{\dot{a}}{a}$ is the Hubble parameter and $\tau=\int\frac{dt}{a}=-\frac{1}{aH}$ is the conformal time for $a=e^{Ht}$. If one has to place quantum theory in dS spacetime Eq. 8, the first and foremost things to be understood are the parity and time reversal operations. The usual conception of understanding the expansion of the Universe is to restrict ourselves to $\tau<0$. But the metric carries a symmetry
@@
Restricting to $\tau<0$ before quantization of a classical field (as is usually followed in standard QFTCS) means throwing away the symmetry in Eq. 9 by hand. The expanding Universe can be described in two ways:
@@
Notice in particular that in Eq. 10, reversing the arrow of time together with the sign of the Hubble parameter represents the same expanding universe. This means that the scale factor is the clock that determines the expansion of the universe. This would explain why in the Wheeler-de Witt equation the cosmic time does not explicitly appear; rather, the scale factor is the clock that runs with the expanding universe Kiefer (2007). Note that $H\to-H$ does not change any properties of the dS space at all. For example, curvature invariants such as the Ricci scalar $R=12H^{2}$ remain the same under a discrete sign flip of H. Eq. 10 illustrates two possible time realizations for an expanding universe. By convention, choosing $\tau>0$ or $\tau<0$ results in loss of information beyond the horizon. In his famous monograph of 1956, Schrodinger Erwin Schrodinger (1956) rejected the idea of two universes because it would lead to a conundrum of ignorance of information beyond the horizon, which in modern language means pure states evolving into mixed states which leads the violation of unitarity and information-loss (Gibbons and Hawking (1977); Parikh et al. (2003)).
Similar to the Minkowski space (see Appendix. B), the quantum fields in the dS space are described through a direct-sum vacuum $|0\rangle_{\rm dS}=|0_{+}\rangle_{\rm dS}\oplus|0_{-}\rangle_{\rm dS}$ corresponding to the direct-sum Fock space $\mathcal{F}_{\rm dS}=\mathcal{F}_{\rm dS+}\oplus\mathcal{F}_{\rm dS-}$. This implies that everywhere in dS space, a quantum field is tied by time forward and backward evolutions at the parity conjugate points, see Appendix. C. At any moment of dS expansion, the parity-conjugate points on the comoving horizon $r_{H}=|\frac{1}{aH}|$ are space-like separated for an imaginary classical observer as shown in Fig. 1.
@@
Note that by DQFT construction, the direct-sum quantum states $|\phi_{\pm}\rangle=\hat{d}_{(\pm)\,\textbf{k}}^{\dagger}|0_{\pm}\rangle_{\rm dS}$ that follow from Eq. 55 are the opposite time evolving positive energy states at the parity conjugate points and this holds for any imaginary classical observer. Furthermore, in DQFT, causality and locality are preserved too. A consequence of this is that any state that disappears beyond the horizon (say at $\left(\theta,\,\varphi\right)$) happens to be the complementary direct-sum (component) state that reappears at the antipodal point (at $\left(\pi-\theta,\,\pi+\varphi\right)$). This is because all the imaginary classical observers, including those on the horizons of each other, should observe the same direct-sum quantum state $|\phi_{+}\rangle\oplus|\phi_{-}\rangle$.
@@
where $|\phi_{m}\rangle$ are the eigenvectors, i.e. for any observable (an operator) $\hat{O}$ we have: $\hat{O}|\phi_{m}\rangle=\lambda_{m}|\phi_{m}\rangle$ where $\lambda_{m}$ are the corresponding eigenvalues. Each $\lambda_{m}$ can be observed in $|\psi\rangle$ with a probability: $P(\lambda_{m})=|d_{m}|^{2}$. Consider now the (de Sitter) Hubble horizon $r_{H}=|\frac{1}{aH}|$ around an observer (as illustrated in Fig. 1). In the standard approach, information is lost (a pure state becomes a mix state) because communication can not happen for space-like distances (i.e. outside the horizon) as was shown by Gibbons and Hawking in 1977 Gibbons and Hawking (1977), in analogy to what happens in Hawking radiation and the black hole information loss paradox Hawking (1974); Kumar and Marto (2024a). This also implies the loss of unitarity in QFT on spacetimes with horizons. In the DQFT approach, we have:
@@
where $|\phi_{Am}^{\pm}\rangle$ and $|\phi_{Bn}^{\pm}\rangle$ are the eigenvectors corresponding to each $\mathcal{P}\mathcal{T}$ separated Hilbert spaces, which remain within the horizon because no information is exchanged across $\mathcal{P}\mathcal{T}$ separated spaces. The fact that there are no cross terms in Eq. 15 follows from the definition of superselection sector Hilbert spaces, according to which one cannot form any superposition of states from the superselection sectors (See Conway (2010); Giulini (2007); nLab authors (2023) for more details on superselection sectors and direct-sum Hilbert spaces and the associated tensor products). This is why in Eq. 15 we do not have any cross-products between parity conjugate states. All information about the entanglement pairs is mapped within the horizon; thus, the pure state will remain pure. In other words, DQFT forbids any entanglement beyond the gravitational horizons. DQFT constructs a direct-sum Hilbert space with the interior and exterior separated by a geometric superselection rule (which is similar to the concept of superselection sectors in QFT Wick et al. (1952); nLab authors (2023)). Thus, there would be an observer complementarity (see Fig. 3 in Kumar and Marto (2025)). We can also understand the same from the point of view of dS space in static coordinates, which reads as